Determine the maximum and minimum of the function f (x) = -x3 + 2x2 – x [0.2].
1
Expert's answer
2019-12-12T12:22:21-0500

"1) We \, find \, f'(x):\\ 
f'(x) = -3x^2 + 4x-1; \\ 
2) f'(x) = 0 \\ 
-3x^2 + 4x - 1 = 0; \\ 
3x^2 - 4x +1 = 0; \\ 
D = b^2 - 4ac = (-4)^2 - 4*3*1 = 16 - 12 = 4; \\ 
x_1 = \frac{-b - \sqrt{D}}{2a} = \frac{4 - 2}{2*3} = \frac{2}{6} = \frac{1}{3}; \\
x_2 = \frac{-b + \sqrt{D}}{2a} = \frac{4 + 2}{2*3} = \frac{6}{6} = 1;"


3)

"4) \newline
f(0) = -0^3+2*0^2-0= 0; \newline
f(\frac{1}{3}) = -(\frac{1}{3})^3 + 2(\frac{1}{3})^2 - \frac{1}{3} = -\frac{4}{27};\newline
f(1) = -1^3 +2*1^2-1 = 0; \newline
f(2) = -2^3 + 2*2^2 - 2 = -2;\\
max \, f=max \{0; -\frac{4}{27};-2\}=0;\\
min \,f=min \{0; -\frac{4}{27};-2\}=-2.\\"

5) Therefore, the function  "f(x) = -x^3 + 2x^2 - x" in the interval "[0,2]"  has its own minimum equal to "-2" (the function attains this minimum at the point "x=2") and the function has the maximum equal to 0 (the function attains this maximum at the points "x = 0" and "x=1").